OpenStudy (anonymous):

The graph of which equation has a vertex of (1, –3)? answer please!!!!!!!!!!!

OpenStudy (mertsj):

You must post some choices.

OpenStudy (anonymous):

y = 1(x – 3)2 + 1 y = 2(x – 3)2 + 1 y = 2(x – 1)2 – 3 y = 2(x + 1)2 – 3

OpenStudy (mertsj):

Third one